### v1

@@ 4.2 Future AI regulation directions
Rapid AI technological development vs slow and robust regulatory frameworks: With the widespread application of AI, a comprehensive regulatory framework is essential to govern AI development and deployment, encompassing various laws related to the environment and labor relations. This survey focuses on ethical issues such as safety, transparency, traceability, non-discrimination, and environmental sustainability. However, the rapid advancement of AI results in continuous breakthroughs and widespread deployment across multiple domains, creating a regulatory lag and involving a complex array of stakeholders, as shown in Figure  3. These stakeholders include enterprises, researchers and developers, regulators and policymakers, end-users, and customers. AI technologies enable rapid innovation cycles with frequent updates and iterations to enhance performance and capabilities, which can sometimes surpass researchers’ and developers’ ability to fully understand and address potential ethical concerns [288, 289]. For example, jailbreaking prompts have recently emerged as an effective method that may missed by end-users and customers to bypass security restrictions and generate harmful content that was originally intended to be prohibited [290].
Required technical complexity and global standardization: Robust regulations and legislative acts necessitate a consensus on values and moral perspectives, as well as a deep understanding of technical complexities to achieve AI goals. This complexity stems from the need to tackle a diverse array of issues while ensuring that AI technologies are safe, transparent, traceable, non-discriminatory, and environmentally sustainable. As discussed in the previous section  4.1, it is crucial to identify technical challenges and trade-off solutions to comprehensively address ethical AI goals. Additionally, given the rapid misuse and abuse of LLMs [288], it is essential to understand related technologies to identify associated risks and corresponding solutions. Our survey provides a comprehensive analysis of AI technologies regarding various ethical issues.

### v2

@@ 4.2 Future AI regulation directions
Rapid AI technological development vs slow and robust regulatory frameworks: With the widespread application of AI, a comprehensive regulatory framework is essential to govern AI development and deployment, encompassing various laws related to the environment and labor relations. This survey focuses on ethical issues such as safety, transparency, traceability, non-discrimination, and environmental sustainability. However, the rapid advancement of AI results in continuous breakthroughs and widespread deployment across multiple domains, creating a regulatory lag and involving a complex array of stakeholders, as shown in Figure  3. These stakeholders include enterprises, researchers and developers, regulators and policymakers, end-users, and customers. AI technologies enable rapid innovation cycles with frequent updates and iterations to enhance performance and capabilities, which can sometimes surpass researchers’ and developers’ ability to fully understand and address potential ethical concerns (?, ?). For example, jailbreaking prompts have recently emerged as an effective method that may missed by end-users and customers to bypass security restrictions and generate harmful content that was originally intended to be prohibited (?).
Required technical complexity and global standardization: Robust regulations and legislative acts necessitate a consensus on values and moral perspectives, as well as a deep understanding of technical complexities to achieve AI goals. This complexity stems from the need to tackle a diverse array of issues while ensuring that AI technologies are safe, transparent, traceable, non-discriminatory, and environmentally sustainable. As discussed in the previous section  4.1, it is crucial to identify technical challenges and trade-off solutions to comprehensively address ethical AI goals. Additionally, given the rapid misuse and abuse of LLMs (?), it is essential to understand related technologies to identify associated risks and corresponding solutions. Our survey provides a comprehensive analysis of AI technologies regarding various ethical issues.